OpenStudy (anonymous):

redict the next two numbers in this sequence: 1/4, 1/9, 1/16, 1/25,...

OpenStudy (anonymous):

1/4 = (1/2)^2 1/9 = (1/3)^2 1/16 = (1/4)^2 1/25 = (1/5)^2

OpenStudy (anonymous):

do you see the sequence?

OpenStudy (anonymous):

1/36,1/49